# Agentar-Scale-SQL: Advancing Text-to-SQL through Orchestrated Test-Time Scaling

### v3

@@ 4.1 Experimental Setup
Implementation Details. We implement Agentar-Scale-SQL with LangChain222https://github.com/langchain-ai/langchain/LangGraph333https://github.com/langchain-ai/langgraph and chroma444https://github.com/chroma-core/chroma retrieval using all-MiniLM-L6-v2555https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 embeddings. The Task Understanding is powered by Gemini-2.5-Flash (Comanici et al., 2025) (temperature 0.2). The ICL SQL Generator utilizes Gemini-2.5-Pro (Comanici et al., 2025) (abbreviated as pro) with two temperature settings (0.5 and 1.8) and GPT-5 (OpenAI, 2025) (temperature 1.0). The Reasoning SQL Generator is fine-tuned based on Omni-SQL-32B (Li et al., 2025b). By default, candidates comprise 9+ from the ICL SQL Generator and 8 from the Reasoning SQL Generator. The SQL Fixer and SQL Reviser both use pro. The base model of the Reasoning SQL Selector is Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
@@ 4.4 Analysis of Generator Components (RQ3)
As shown in Figure 3, the ICL generator achieves a notably higher upper bound accuracy (81.36%) than the reasoning generator (75.88%), indicating its strong potential for generating correct queries. However, combining their outputs (All) achieves the highest overall upper bound of 84.29%. This synergistic gain is explained by their complementary nature, as illustrated in Figure 4. While they share a large set of correct solutions, they also uniquely solve 47 and 12 samples, respectively. This expanded coverage holds across all difficulty levels (Figure 5). A breakdown by difficulty reveals that the reasoning generator holds an edge on simple and moderate tasks, while the ICL generator proves more effective for solving challenging problems.
@@ 5 Conclusion
Building on the success of enhancing intelligence through Test-Time Scaling, we are now pioneering our next great endeavor: Exercise-Time Scaling. We will empower a new generation of agents to learn through action and evolve from experience, on a journey to realize intelligence that transcends humanity and to explore the vast unknown.
@@ Appendix A BIRD Leaderboard: Bar Chart and Snapshot
We present the BIRD rankings in a bar chart (Figure 7) and provide a snapshot of the BIRD leaderboard (Figure 8).
@@ Appendix D Discussion of Schema Linking
Appendix E Further Work

### v4

@@ 4.1 Experimental Setup
Implementation Details. We implement Agentar-Scale-SQL with LangChain222https://github.com/langchain-ai/langchain/LangGraph333https://github.com/langchain-ai/langgraph and chroma444https://github.com/chroma-core/chroma retrieval using all-MiniLM-L6-v2555https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 embeddings. The Task Understanding is powered by Gemini-2.5-Flash (Comanici et al., 2025) (temperature 0.2). The ICL SQL Generator utilizes Gemini-2.5-Pro (Comanici et al., 2025) (abbreviated as pro) with two temperature settings (0.5 and 1.8) and GPT-5 (OpenAI, 2025) (minimal reasoning effort). The Reasoning SQL Generator is fine-tuned based on Omni-SQL-32B (Li et al., 2025b). By default, candidates comprise 9 from the ICL SQL Generator and 8 from the Reasoning SQL Generator. The SQL Fixer and SQL Reviser both use pro. The base model of the Reasoning SQL Selector is Qwen2.5-Coder-32B-Instruct (Hui et al., 2024). The RL framework leverages verl (Sheng et al., 2025) on 32 NVIDIA A100 80GB GPUs.
@@ 4.4 Analysis of Generator Components (RQ3)
As shown in Figure 3, the ICL generator achieves a notably higher upper bound accuracy (81.36%) than the reasoning generator (75.88%), indicating its strong potential for generating correct queries. However, combining their outputs (All) achieves the highest overall upper bound of 84.29%. This synergistic gain is explained by their complementary nature, as illustrated in Figure 4. While they share a large set of correct solutions, they also uniquely solve 47 and 12 samples, respectively. This expanded coverage holds across all difficulty levels (Figure 5). A breakdown by difficulty reveals that the reasoning generator holds an edge on simple and moderate tasks, while the ICL generator proves more effective in challenging problems.
@@ 5 Conclusion
Building on the success of enhancing intelligence through Test-Time Scaling, we are pioneering our next endeavor: Exercise-Time Scaling. We will empower a new generation of agents to learn through action and evolve from experience.
@@ Appendix A BIRD Leaderboard: Bar Chart and Snapshot
We present the BIRD rankings as a bar chart (Figure 7) and a snapshot (Figure 8) as of September 28, 2025.
@@ Appendix D Discussion of Schema Linking
Appendix E Prompts
We have listed a selection of prompts. For the complete list, please refer to the code repository.
The prompt for the Task Understanding.
The prompt for the Reasoning Generator.
The prompt for the ICL Generator with CoT prompting.
The prompt for the SQL fixer.
The prompt for the Reasoning Selector.
Appendix F Further Work